How do you factor m with a power of 2, -24m -81.
Answer by oberobic(2304)

m^2 -24m -81 = 0
.
You need to look at factors of 81 that are 24 apart.
Otherwise you will have to use the quadratic equation.
.
3 * 27 = 81
and
27 - 3 = 24
so
you can factor it
.
(m - 27)(m + 3) = 0
.
Use FOIL to check
.
m^2 +3m -27m -81 = 0
.
m^2 -24m -81 = 0
.
Done.